# Conditionally valid Probabilistic Conformal Prediction

### v1

@@ 3 Theoretical guarantees
In this section, we provide both marginal and conditional guarantees for the prediction set $\mathcal{C}_{\alpha}(x)$ given in (4). The validity of these guarantees is ensured by the exchangeability of the calibration data, with the exception of Corollary 3.3 which relies on a concentration inequality and thus requires that the calibration data are i.i.d. The following theorem establishes marginal validity of the predictive set defined by CP2.
@@
Assume 1-2. Then, for any $\alpha\in(0,1)$, it holds
$1-\alpha\leq\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right).$
Moreover, if the conformity scores $\{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})\}_{k=1}^{n+1}$ are almost surely distinct, then it also holds that
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)<1-\alpha+\frac{1}{n+1}.$
The proof is postponed to Section A.1. Moreover, the upper bound on the coverage always holds when the distribution of $f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})$ is continuous. Now, we will investigate the conditional validity of $\mathcal{C}_{\alpha}(x)$. Denote by $\mathrm{d}_{\mathrm{TV}}$ the total variation distance.
@@
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(x)\mid X_{n+1}=x,Z_{n+1}=z}\right)\geq 1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})-p_{n+1}^{(x,z)},$
@@
where $p_{n+1}^{(x,z)}=\mathbb{P}\left({Q_{1-\alpha}(\mu)<f_{\bar{\tau}_{n+1}}^{-1}(\bar{\lambda}_{n+1})\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right)$.
The proof is postponed to Section A.1. The better the estimator $\Pi_{Y\mid X=x}$ is, the closer the result is to $1-\alpha$. Achieving accurate conditional coverage at $x$ does not require knowledge of the entire conditional distribution $\textup{P}_{Y\mid X}$. Instead, only a reliable approximation for the specific point $x$ is required. dThe second term in the lower bound is $p_{n+1}^{(x,z)}$. Its expected value is upper bounded by $\mathbb{E}[p_{n+1}^{(x,z)}]\leq\alpha$, but analyzing this term is difficult. We address this problem in the following corollary, whose proof is postponed to Section A.2. Denote by $F$ and $\hat{F}$ the cumulative distribution functions of the random variables $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ and $f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})$, where $(X,Y,Z)\sim\textup{P}_{X}\otimes\textup{P}_{Y\mid X}\otimes\Pi_{Z\mid X}$ and $(X,\hat{Y},Z)\sim\textup{P}_{X}\otimes\Pi_{Y\mid X}\otimes\Pi_{Z\mid X}$, respectively.
Corollary 3.3.
Assume 1-2, and suppose the distribution of $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ is continuous. For any $x\in\mathbb{R}^{d}$, $z\in\mathcal{Z}$ and $\epsilon\in[0,1-\alpha)$, it holds that
$p_{n+1}^{(x,z)}\leq\exp\left({-n\Phi(\epsilon)}\right)+\mathbb{P}\left({F^{-1}(1-\alpha-\epsilon)<f_{\bar{\tau}_{n+1}}^{-1}(\bar{\lambda}_{n+1})\leq\hat{F}^{-1}(1-\alpha)\,\big{|}\,X_{n+1}=x,Z_{n+1}=z}\right)$
where $\Phi(\epsilon)=\epsilon[(u_{\epsilon}^{-1}-1)\log(1+u_{\epsilon})-1]$ and $u_{\epsilon}=\epsilon(\alpha+\epsilon)^{-1}(1-\alpha-\epsilon)^{-1}$.
Setting $\epsilon=\sqrt{8\alpha(1-\alpha)n^{-1}\log n}$ ensures that $\exp(-n\Phi(\epsilon))\leq n^{-1}$. When $F^{-1}$ is continuous, $F^{-1}(1-\alpha)$ is approximately equal to $F^{-1}(1-\alpha-\epsilon)$ for small $\epsilon$. Therefore, if $\hat{F}^{-1}(1-\alpha)$ closely approximates $F^{-1}(1-\alpha)$, then Corollary 3.3 demonstrates that $p_{n+1}^{(x,z)}$ is always small.
[…]
$\bar{\mathcal{C}}_{\alpha}(x)=\mathcal{R}_{z}\left({x;f_{\tau_{x,z}}\Big{(}{Q_{(1-\alpha)(1+n^{-1})}\big{(}{\textstyle\frac{1}{n}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}}\big{)}}\Big{)}}\right).$
Corollary 3.4.
The proof of Corollary 3.4 is along the same lines as [41, Corollary 1]; see Section A.3.
@@ 4 Numerical experiments
In this section, we conduct a comprehensive analysis demonstrating the advantage of CP2 compared to standard and adaptive split conformal algorithms. Specifically, we benchmark our algorithm against two state-of-the-art methods: Conformalized Quantile Regression (CQR; [30]) and Probabilistic Conformal Prediction (PCP; [41]). We aim to answer these specific questions: how does CP2 performs in terms of coverage, conditional coverage and predictive set volume when compared to state-of-the-art methods on synthetic and real data.
4.1 Synthetic data experiments
In this example, $(X_{k},Y_{k})$ is sampled from a mixture of $P=4$ Gaussians; see Figure 4(a). The results for other classical 2-d datasets lead to similar conclusions. The number of training and calibration samples is $T=10^{4}$ and $n=10^{3}$, respectively. We fit a Mixture Density Network (MDN) as an explicit generative model, $\gamma_{Y\mid X=x}(y)=\sum_{\ell=1}^{P}\pi_{\ell}(x)\mathcal{N}(y;\mu_{\ell}(x),\sigma^{2}_{\ell}(x))$, where $\mu_{\ell}(\cdot)$, $\sigma_{\ell}(\cdot)$ and $\pi_{\ell}(\cdot)$ are all modeled by fully connected 2-layers neural networks (the condition $\sum_{\ell=1}^{P}\pi_{\ell}(x)=1$ is ensured by using softmax activation functions). We use CP2-HPD (the calculation of the HPD rates as well as $\tau_{x}$ and $\lambda_{x,y}$ is explicit in this case). The parameters of the MDN are trained by maximizing the likelihood on the training set.
We compare the plain CP2-HPD, PCP (with the same MDN as CP2-HPD and $M=50$ draws) and CQR. All methods achieve the desired nominal coverage $0.91-\alpha=$ 0.9. We illustrate the conditional coverage in Figure 4(b) and the lengths of the predictive sets in Figure 4(c). CP2-HPD with a fixed-width predictive set performs poorly in this multimodal example, both in terms of the size of the confidence set and the conditional coverage. CP2-HPD and CQR perform similarly in terms of conditional coverage (which remains close to $0.91-\alpha=$ 0.9). The conditional coverage of PCP varies between 0.85 and 0.95. CP2 produces shorter prediction sets compared to CQR and PCP. This is because CP2 uses an HPD confidence set that is more suitable for multimodal applications than the confidence set used by CQR.
@@ 4.1 Synthetic data experiments
4.2 Real data experiments
In this section, we study the performance of CP2-PCP on several real world regression datasets.
Datasets. We use publicly available regression datasets, which are also considered in [30, 41]. Some of them come from the UCI repository: bike sharing (bike), protein structure (bio), blog feedback (blog), Facebook comments (fb1 and fb2). Other datasets come from US Department of Health surveys (meps19, meps20 and meps21), and from weather forecasts (temp[10].
Methods. We compare the proposed CP2-PCP method with Probabilistic Conformal Prediction (PCP; [41]), Conformalized Quantile Regression (CQR; [30]) and Conformalized Histogram Regression (CHR; [35]). We also consider CQR2222 which is a modification of CQR that uses inverse quantile nonconformity score. For our method and PCP we use a Mixture Density Network [4] to estimate the conditional distribution $\textup{P}_{Y\mid X}$, since it was chosen in [41] as best-performing. We also consider different choices of $f_{\tau}$ for our method: CP2-PCP-L stands for CP2-PCP with $f_{\tau}(\lambda)=\lambda\tau$ and CP2-PCP-D stands for CP2-PCP with $f_{\tau}(\lambda)=\lambda+\tau$. Our implementation of CP2-PCP is summarized in Algorithms 1 and 2. They use Brent’s method to find the optimal $\tau_{x}$; see (3). Additionally, we consider $\Pi_{Y\mid X}$ which is a special case of CP2-PCP with $f_{\tau}(\lambda)=\tau$.
Metrics. Empirical coverage (marginal and conditional) is the main quantity of interest for prediction sets. We evaluate worst-slab conditional coverage [8, 31] in our experiments, see details in Appendix B.3. We also measure the total size of the predicted sets, scaled by the standard deviation of the response $Y$.
Experimental setup. Our experimental setup closely follows [41]. We split each dataset into train, calibration and test portions and train a Mixture Density Network with $10$ components to approximate $\textup{P}_{Y\mid X}$. For each calibration and test point we first obtain the Gaussian Mixture parameters (this becomes $\Pi_{Y\mid X}$) and then sample $M=5,20,50$ samples from these distributions (gives us $\mathcal{R}_{z}(x,\tau)$). We replicate this experiment on $50$ different splits of each dataset.
Results of the experiments for $M=50$ samples are presented in Figures 5 and 6, additional results are available in Appendix B. In terms of marginal coverage, all methods achieve the target $1-\alpha$ value, except for $\Pi_{Y\mid X}$.
@@ 4.2 Real data experiments
Additionally, we assess conditional coverage with the help of clustering. We apply HDBSCAN [7, 26] method to cluster the test set and then compute coverage within clusters. Results for fb1 dataset are presented in Figure 7. We again observe that CP and PCP do not achieve conditional coverage and CHR and CQR performance is unstable. CP2-PCP on the other hand maintains valid conditional coverage on all clusters and even on outliers (cluster label -1). Not that these are all outliers combined and they may not lie in the same region of the input space.

### v2

@@ 3 Theoretical Guarantees
In this section, we provide both marginal and conditional guarantees for the prediction set $\mathcal{C}_{\alpha}(x)$ given in (4). The validity of these guarantees is ensured by the exchangeability of the calibration data, with the exception of Theorem 3.3 which relies on a concentration inequality and thus requires i.i.d. calibration data. The following theorem establishes marginal validity of the predictive set defined by CP2.
@@
Assume 1-2. Then, for any $\alpha\in(0,1)$, it holds $1-\alpha\leq\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)$. Moreover, if the conformity scores $\{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})\}_{k=1}^{n+1}$ are almost surely distinct, then it also holds that $\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)<1-\alpha+(n+1)^{-1}.$
The proof is postponed to Section A.1. Moreover, the upper bound on the coverage always holds when the distribution of $f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})$ is continuous. Now, we will investigate the conditional validity. Denote by $\mathrm{d}_{\mathrm{TV}}$ the total variation distance and by $\mathbb{P}^{\mathcal{T}}$ the conditional probability given the training data.
@@
$\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(x)\mid(X_{n+1},Z_{n+1})=(x,z)}\right)\geq 1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})-p_{n+1}^{(x,z)},$
@@
where $p_{n+1}^{(x,z)}=\mathbb{P}^{\mathcal{T}}\left({Q_{1-\alpha}(\mu_{n})<f_{\bar{\tau}_{n+1}}^{-1}(\bar{\lambda}_{n+1})\leq\varphi\,|\,(X_{n+1},Z_{n+1})=(x,z)}\right)$.
The proof is postponed to Section A.1. The more accurately the estimator $\Pi_{Y\mid X=x}$ approximates the true conditional distribution, the closer the result will be to $1-\alpha$. The second term in the lower bound is $p_{n+1}^{(x,z)}$. Its expected value is upper bounded by $\mathbb{E}[p_{n+1}^{(X,Z)}]\leq\alpha$, and non-asymptotic bounds for this error term are developed in Section A.2.
We will now briefly discuss the asymptotic conditional coverage guarantee; details are provided in the supplementary paper. Assuming the availability of an oracle for the predictive distribution, i.e., $\textup{P}_{Y\mid X=x}=\Pi_{Y\mid X=x}$,we get under 1 and 2, that for any $t\in\mathbb{R}$,
$=\Pi_{Y\mid X=x}\left({\mathcal{R}_{z}(x;f_{\tau_{x,z}}(t))}\right),$
where $(X,Y,Z)$ follows the same distribution than $(X_{k},Y_{k},Z_{k})$, $k\in\{1,\dots,n\}$. Note that $\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(t)))\geq 1-\alpha$ if and only if $t\geq\varphi$, which implies that
$\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})\leq t\mid(X,Z)=(x,z))\geq 1-\alpha$   if and only if   $t\geq\varphi$. (6)
From (6) it is easily seen that the $(1-\alpha)$-quantile of $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ is $\varphi$. The Glivenko–Cantelli Theorem (Van der Vaart, 2000, Theorem 19.1) demonstrates that $\sup_{t\in\mathbb{R}}|\mu_{n}(-\infty,t]-\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})\leq t)|\to 0$ almost surely as $n\to\infty$, where $\mu_{n}$ is defined in (5). Since the convergence of the c.d.f. implies the convergence of the quantile function (Van der Vaart, 2000, Lemma 21.2), we deduce that $Q_{1-\alpha}(\mu_{n})\to\varphi$ almost-surely as $n\to\infty$. Under weak additional conditions this implies that $\lim_{n\to\infty}p_{n+1}^{(x,z)}=0$, $\bar{\Pi}_{Z\mid X}\times\textup{P}_{X}$-almost everywhere, where $\bar{\Pi}_{Z\mid X}$ is the Markov kernel used to draw the auxiliary variables $z$; see Section A.4. In this case, Theorem 3.1 implies the asymptotic validity of CP2.
In practice, the oracle is unavailable. In the following theorem, we examine the asymptotic conditional conformal validity as the size of the training dataset, $m_{n}$, goes to infinity with $n$. In most cases, $\lim_{n\to\infty}m_{n}/n=\gamma>0$, but this is not required here. To make the dependency of the estimator on the size of the training set explicit, we will denote the conditional distribution a $\Pi_{Y\mid X}^{(m_{n})}$. Consider the following assumption.
H 3.
There exists sequence $(r_{n})$ such that $\lim\limits_{n\to\infty}\mathbb{P}(\mathrm{d}_{\mathrm{TV}}(\textup{P}_{X,Y};\textup{P}_{X}\times\Pi_{Y\mid X}^{(m_{n})})\leq r_{n})=1.$
In most interesting case, we have $\lim_{n\to\infty}r_{n}=0$. Such types of bounds can be deduced from Devroye & Lugosi (2001, Chapter 9). Let $(X,Y,Z)$ and $(X,\hat{Y},Z)$ be random variables distributed according to $\textup{P}_{X,Y}\times\bar{\Pi}_{Z\mid X}$ and $\textup{P}_{X}\times\Pi_{Y\mid X}^{(m_{n})}\times\bar{\Pi}_{Z\mid X}$, respectively.
Theorem 3.3.
Assume 1-2-3 hold. If the distributions of $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ and $f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})$ are continuous, then, it holds
$\left\lvert{\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\,|\,X_{n+1},Z_{n+1}}\right)-1+\alpha}\right\rvert=\operatorname{\mathrm{O}}_{\mathbb{P}}\left({\sqrt{n^{-1}\log n}+r_{n}}\right).$
In Lei et al. (2018); Izbicki et al. (2020); Sesia & Candès (2020), the asymptotic conditional validity is demonstrated by assuming the consistency of their methods’ estimators. For instance, Romano et al. (2019) assume that the conditional quantile regressor converges in $L^{2}$ towards the true quantile with high probability.
@@ 4 Numerical Experiments
In this section, we conduct a comprehensive analysis demonstrating the advantage of CP2 compared to standard and adaptive split conformal algorithms. Specifically, we benchmark our algorithm against several state-of-the-art methods: Conformalized Quantile Regression (Romano et al., 2019), Conformalized Histogram Regression (Sesia & Romano, 2021) and Probabilistic Conformal Prediction (Wang et al., 2023). All these method share some key aspects: they are built on top of the pre-trained models and do not require access to training data or the model’s internals on both calibration and prediction steps. We aim to answer these specific questions: how does CP2 performs in terms of coverage, conditional coverage and predictive set volume when compared to state-of-the-art methods on synthetic and real data.
4.1 Synthetic Data Experiment
In this example, $(X_{k},Y_{k})$ is sampled from a mixture of $P=4$ Gaussians; see Figure 2(a). The number of training and calibration samples is $m=10^{4}$ and $n=10^{3}$, respectively. We fit a Mixture Density Network (MDN) as an explicit generative model, $\gamma_{Y\mid X=x}(y)=\sum_{\ell=1}^{P}\pi_{\ell}(x)\mathcal{N}(y;\mu_{\ell}(x),\sigma^{2}_{\ell}(x))$, where $\mu_{\ell}(\cdot)$, $\sigma_{\ell}(\cdot)$ and $\pi_{\ell}(\cdot)$ are all modeled by fully connected 2-layers neural networks (the condition $\sum_{\ell=1}^{P}\pi_{\ell}(x)=1$ is ensured by using softmax activation functions). We use CP2-HPD (the calculation of the HPD rates as well as $\tau_{x}$ and $\lambda_{x,y}$ is explicit in this case) with $f_{t}(v)=tv$. The parameters of the MDN are trained by maximizing the likelihood on the training set.
We compare the plain CP2-HPD, PCP (with the same MDN as CP2-HPD and $M=50$ draws) and CQR. All methods achieve the desired marginal coverage $0.91-\alpha=$ 0.9. We illustrate the conditional coverage in Figure 2(b) and the lengths of the predictive sets in Figure 2(c). CP with a fixed-width predictive set performs poorly in this multimodal example, both in terms of the size of the confidence set and the conditional coverage. CP2-HPD and CQR perform similarly in terms of conditional coverage (which remains close to $0.91-\alpha=$ 0.9). The conditional coverage of PCP varies between 0.85 and 0.95. CP2-HPD produces shorter prediction sets compared to CQR and PCP. This is because CP2-HPD uses an HPD confidence set that is more suitable for multimodal applications than the interval produced by CQR.
@@ 4.1 Synthetic Data Experiment
4.2 Real-world Regression Data Experiments
In this section, we study the performance of CP2-PCP on several real world regression datasets. In particular, we used the same datasets as in Wang et al. (2023) to facilitate comparisons.
Datasets. We use publicly available regression datasets, which are also considered in Romano et al. (2019); Wang et al. (2023). Some of them come from the UCI repository: bike sharing (bike), protein structure (bio), blog feedback (blog), Facebook comments (fb1 and fb2). Other datasets come from US Department of Health surveys (meps19, meps20 and meps21), and from weather forecasts (temp(Cho et al., 2020).
Methods. We compare the proposed CP2-PCP method with Probabilistic Conformal Prediction (PCP; Wang et al. (2023)), Conformalized Quantile Regression (CQR; Romano et al. (2019)) and Conformalized Histogram Regression (CHR; Sesia & Romano (2021)). We also consider CQR2222 which is a modification of CQR that uses inverse quantile nonconformity score. For our method and PCP we use a Mixture Density Network Bishop (1994) to estimate the conditional distribution $\textup{P}_{Y\mid X}$, since it was chosen in Wang et al. (2023) as best-performing. We also consider different choices of $f_{t}$ for our method: CP2-PCP-L stands for CP2-PCP with $f_{t}(v)=tv$ and CP2-PCP-D stands for CP2-PCP with $f_{t}(v)=t+v$. Our implementation of CP2-PCP is summarized in Algorithm 1. Additionally, we consider $\Pi_{Y\mid X}$ which is a special case of CP2-PCP with $f_{t}(v)=t$.
Metrics. Empirical coverage (marginal and conditional) is the main quantity of interest for prediction sets. We evaluate worst-slab conditional coverage (Cauchois et al., 2020; Romano et al., 2020b) in our experiments, see details in Appendix B.3. We also measure the total size of the predicted sets, scaled by the standard deviation of the response $Y$.
Experimental setup. Our experimental setup largely follows the approach outlined in Wang et al. (2023). Specifically, we split each dataset into training, calibration, and testing sets. A Mixture Density Network (MDN) with 10 components is then trained to approximate the conditional distribution $\textup{P}_{Y\mid X}$. For each calibration and test point, we first compute the Gaussian Mixture parameters, forming $\Pi_{Y\mid X}$, and subsequently draw $M=5,20,50$ samples from these distributions, which yield $\mathcal{R}_{z}(x,t)$. This process is repeated across 50 different random splits of each dataset. Results of the experiments for $M=50$ samples are presented in Figures 3 and 4, additional results are available in Appendix B. In terms of marginal coverage, all methods achieve the target $1-\alpha$ value, except for $\Pi_{Y\mid X}$.
@@ 4.2 Real-world Regression Data Experiments
Finally, we assess conditional coverage with the help of clustering. We apply HDBSCAN (Campello et al., 2013; McInnes & Healy, 2017) method to cluster the test set and then compute coverage within clusters. Results for fb1 dataset are presented in Figure 5. We again observe that CP and PCP do not achieve conditional coverage and CHR and CQR performance is unstable. CP2-PCP on the other hand maintains valid conditional coverage on all clusters and even on outliers (cluster label -1). Note that these are all outliers combined and they may not lie in the same region of the input space.
@@
Figure 6: Marginal coverage for multi-target datasets, 50 replications. Sample size was set to 1000. Nominal coverage equals $(1-\alpha)=0.9$ and is shown by dashed black line.
4.3 Real-world Regression Data with Multi-dimensional Targets
We also study CP2 family of algorithms on the multi-target regression problems. Since selecting the threshold $\tau$ for our methods is not dependent on the number of dimensions in $Y$ their application is straightforward. On the other hand, most other methods are inherently one-dimensional thus require the use of the Bonferroni correction (Dunn, 1961). Each coordinate is treated independently with miscoverage level adjusted to $\alpha/d$, where $d$ is the number of targets. As a result, for quantile regression-based methods prediction sets are rectangular cuboids, formed as a product of the corresponding intervals.
Datasets. We consider open-source multidimensional regression datasets: river flow data rf1 and rf2 (Xioufis et al., 2012), supply chain management scm1d and scm20d (Xioufis et al., 2012), indoor localisation indoor (Torres-Sospedra et al., 2014), GPU computation time sgemm_small [^1] The full dataset contains 241600 examples. Due to computational constraints we randomly subsample 10000 examples for each replication of our experiment. (Ballester-Ripoll et al., 2017).
We use the same metrics as before: marginal coverage and worst-slab coverage. Evaluating the difference in prediction set size is more complex in case of multiple dimensions. Due to computational constraints we perform pairwise comparisons between our methods and selected baselines, measuring approximate areas of 2D projections of the prediction sets  (Wang et al., 2023). These results can be found on Figure 8. We approximate areas using a grid and fewer samples.
Since our methods naturally extend beyond one dimension, the experimental setup is almost identical. We use the same underlying model for $\textup{P}_{Y\mid X}$, the prediction set is now a union of $d$-dimensional balls of the same radius around the sampled centers. The number of samples is increased to 1000.
Figure 7: Conditional coverage for multi-target datasets targets, 50 replications. Sample size was set to 1000. Nominal coverage equals $(1-\alpha)=0.9$ and is shown in dashed black. Worst-slab coverage parameter $(1-\delta)=0.1$.
Figure 8: Average rank of the projected set size. For each pair of targets area of the corresponding 2D projection of the prediction set is calculated. For each test point and each pair of targets methods are ranked. Lower rank is smaller area. This graph shows averaged results of 10 replications.
Results. In Figure 6 we show marginal coverage attained by different algorithms. As expected, naive application of 1D techniques CQR, CQR2 and CHR to multiple outputs produces significant overcover. PCP and CP2 methods naturally extend to multidimensional targets and provide correct marginal coverage.
In Figure 7 we present the conditional coverage estimates for multi-target datasets. PCP significantly undercovers on rf1, rf2 and scm1d datasets, while CP2 comes very close to the nominal coverage of $0.9$. In case of CQR, CQR2 and CHR, they still overcover (scm20d, sgemm) or perform comparably to our approach.
Figure 8 shows the aggregated results of the set size comparisons in multidimensional target setting. For each test point and each pair of axes we rank the methods by the area of the projection of the corresponding prediction set. The plot show average rank for each method, aggregated across all axes pairs and replications. Lower rank corresponds to smaller area, which is our goal. For datasets indoor, scm1d and sgemm_small our approach performs better, while also providing sharper conditional covergae, as was shown earlier. On the remaining datasets CP2performs similarly to the competitors.